Abcde is a regular hexagon .prove rhat ab+ac+ad+ae+af=6ao

Grade:11

1 Answers

Deepak Kumar Shringi
5 years ago

AB+AC+AD+AE+AF
=(AB+AE)+(AC+AF)+AD
=AD+AD+AD+=3AD=3(2AO)=6AO
therefore,AB+AC+AD+AE+AF=3(2AD)=6AO.